LASER PHOTOLYSIS OF THE N_{2}O/I_{2}/CO_{2}/Ar MIXTURES AT 193 NM

\underline{M. H. KABIR},V. AZYAZOV, I. O. ANTONOV, and M. C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Oxygen-iodine lasers that utilize electrical discharges to produce O_{2}(a^{1}Δ_{g}) are currently being developed. The discharge generators differ from those used in conventional chemical oxygen iodine lasers in that they produce significant amounts of atomic oxygen. The post-discharge chemistry includes channels that are both beneficial and detrimental to the laser. The beneficial reactions result in the dissociation of I_{2} while the detrimental processes cause direct and indirect removal of I(^{2}P_{1/2}) (the upper level of the laser). We have examined kinetic processes relevant to the laser through studies of photo-initiated reactions in N_{2}O/I_{2} mixtures. The reactions have been monitored using absorption spectroscopy, LIF, and time-resolved emission spectroscopy. A kinetic model has been developed from these data, and the rate constants for the deactivation of I(^{2}P_{1/2}) by O(^{3}P) (k = 1.2 x 10^{−11} cm^{3} s^{−1}) has been determined for the first time.